OpenStudy (anonymous):

given f(x)=5-x/x+5 and g(x)=x+4 find (fog)(x) do I just replace g(x) into fx(x) where the x's are?

OpenStudy (ranga):

(fog)(x) = f(g(x)) Take f(x) = 5-x/x+5 f(g(x)) means everywhere that x occurs, replace it by g(x) or by (x+4)

OpenStudy (anonymous):

1-x/x+9 would that be the correct answer?

OpenStudy (ranga):

Can you write f(x) using the equation editor so it is clearer?

OpenStudy (anonymous):

oh yes sorry

OpenStudy (anonymous):

\[f(x)=\frac{ 5-x }{ x+5 }\]

OpenStudy (anonymous):

would turn into \[f(x)= \frac{ 5-(x+4) }{ (x+4) +5}\]

OpenStudy (ranga):

Okay when writing such expression without the equation editor you should use parenthesis: f(x) = (5 - x) / (x + 5). Yes, your answer for f(g(x)) is correct, except use parenthesis when not using equation editor: f(g(x)) = (1 - x) / (x + 9)

OpenStudy (anonymous):

ohhh okay thank you

OpenStudy (anonymous):

and if im asked to find:g ( f (9) ). I would plug 9 into g(x)?

OpenStudy (ranga):

Look at the innermost function first. Find f(9) by plugging 9 into f(x). Then plug that value into g(x) to get g(f(9))

OpenStudy (anonymous):

ohhh that's why there is )) im going to post another one =)

OpenStudy (ranga):

yes. okay.